### v1

# Semantica: An Adaptable Image-Conditioned Diffusion Model

Abstract
We investigate the task of adapting image generative models to different datasets without finetuneing. To this end, we introduce Semantica, an image-conditioned diffusion model capable of generating images based on the semantics of a conditioning image. Semantica is trained exclusively on web-scale image pairs, that is it receives a random image from a webpage as conditional input and models another random image from the same webpage. Our experiments highlight the expressivity of pretrained image encoders and necessity of semantic-based data filtering in achieving high-quality image generation. Once trained, it can adaptively generate new images from a dataset by simply using images from that dataset as input. We study the transfer properties of Semantica on ImageNet, LSUN Churches, LSUN Bedroom and SUN-397.
1 Introduction
ImageNet
Refer to caption      Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
SUN397
Refer to caption Refer to caption
Refer to caption Refer to caption
LSUN Churches
Refer to caption Refer to caption
LSUN Bedroom
Refer to caption Refer to caption
Figure 1: Left: Conditioning image Right: Five samples at guidance 0.5. Semantica is trained exclusively on web-image pairs. During adaptation, it receives a conditioning image and generates samples reflective of semantic information. Semantica requires no label supervision or finetuning.
Machine learning initially focused on optimizing and improving models on small datasets. The field has transitioned to training general purpose models on web-scale data and then finetuning them for specific tasks on smaller datasets. This paradigm shift has lead to state-of-the-art results on a number of different domains. In this paper, we focus on the relatively underexplored task of adapting an image generative model to different datasets. One approach is to simply train a generative model on a large dataset of unlabelled images and finetune them on smaller datasets. While this approach is straight-forward in theory, it requires clever architecture or regularizer design to prevent overfitting in practice (See Sec.2.1). As models scale up, finetuning for every dataset also just becomes increasingly impractical.
Instead, we apply the principle of in-context learning to adapt a image generative model. We train our model called Semantica using image episodes, which are image pairs that belong to the same web page. Therefore, training relies exclusively on the hypothesis that images from the same web page should have some common semantic attributes. For example, it is highly likely that images scraped from a Wikipedia page on dolphins, contain pictures of dolphins. At adaptation, the model receives an image and then generates another image that preserves semantic information.
Semantica consists of a pre-trained image encoder and a diffusion model. The encoder extracts representations from the conditioning image and guides the diffusion model. We thoroughly investigate the impact of different choices for the image encoder (contrastive or self-supervised) and the conditioning mechanisms (global or local). Our experiments also highlight the importance of semantic data filtering to generate high-quality images. Each row in Fig 1 showcases a conditioning image to the left and five generated samples to the right. Semantica is capable of generating high quality and diverse images, reflective of semantic information from the conditioning image across a number of different datasets.
2 Related work
2.1 Generative Transfer
There has been extensive research that studies the adaptation of source-pretrained generative models to a target dataset with adaptation of weights. Initial works study the transfer of discriminators and generators in GANs from a source dataset to a target dataset [47]. Further, [17] show that ImageNet pretraining on a large GAN model is beneficial for transfer to small datasets. A number of papers focus on improving generation quality by adapting only a subset of parameters. These include scale and shift parameters [32], updating only the higher discriminator layers [28], linear combinations of scale and shift parameters [41], modulating kernels or convolutions [59, 58, 10, 2] and singular values [38], mapping networks from noise to latents [46, 29, 53] and latent offsets [12]. Various works apply regularization losses by enforcing constraints to samples/weights by the source generator including elastic weight regularization [27], domain correspondence [33, 16, 22], contrastive learning [60], spatial alignment [51], inversion [49, 23, 44], random masks on discriminators [61] and alignment free spatial correlation [30]. Given the increasing popularity of VQ-VAE and diffusion based models, recent work [43] and [61] explore few-shot finetuning on VQ-VAE tokens and diffusion models. We defer to the survey [1] for a detailed exposition of all these methods. In contrast to these works, we explore training a generator on web-scale images and study their transfer to standard small-scale image datasets. Retrieval augmented models [7, 5] compute nearest neighbours for a query image across a bank of memory images. These retrieved neighbors facilitate the training or generation process. Unlike these methods, we do not require access to a memory bank of images during train or test time.
2.2 Diffusion
Diffusion and score-based generative models have become increasingly successful in modelling images [20, 40, 31, 4], videos [18, 42] and audio [25]. As generation quality has steadily improved, they have been used in contexts with more and more conditioning variables. Well-known examples are text-to-image and text-to-video modelling, where the conditioning variable is text. In this case, the conditioning variable can be seen as a sequence from which cross-attention layers communicate to the feature maps of the image or video, i.e. what the diffusion model is learning to generate [40, 31]. As the desire for controllable generation increases, solutions such as ControlNet [56] have been developed. ControlNet takes in conditioning images of the same size as the generations, and uses a copy of the UNet to learn an encoder for the conditioning signals. Although this encoder trains fast due to parameter initialization from a pretrained diffusion UNet, it is difficult to deal with different sized inputs. In those cases, only conditioning via cross-attention as done in [52] overcomes the in-place additions between the ControlNet encoder and the base UNet. Conditioning on images as context has produced impressive results, turning scribbles or edge detections into high quality image generations [48] and discriminative tasks [3, 26]. In contrast with the above mentioned techniques, our framework relies on general web-based pretraining for semantic-based adaptive image generation. Finally, while [13] study the transfer of few-shot diffusion models between small datasets (CIFAR-100 -> miniImageNet, we see in Sec. 5.6 that this can lead to sub-optimal transfer.
3 Background
3.1 Diffusion
Diffusion models learn to generate examples by gradually denoising a diffusion process. For a single datapoint, their loss can be expressed as a squared error between the original datapoint and its prediction:
$\mathbb{E}_{t\sim\mathcal{U}(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,{\mathbf{I}})}\big{[}w(t)||{\bm{x}}-f({\bm{z}}_{t})||^{2}\big{]}\text{ where }{\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}{\bm{{\epsilon}}}_{t}$ (1)
It is helpful to define $\mathrm{SNR}(t)=\alpha_{t}^{2}/\sigma_{t}^{2}$. In the case of $w(t)=\mathrm{SNR}(t)$ the loss above is equivalent to a loss in ${\bm{{\epsilon}}}$-space, the simple loss from [19]. After training, the denoising model generates samples by taking small steps. Starting at $t=1$ with initial Gaussian noise and one slowly denoises for timesteps $t=1,1-1/N,\ldots$ where the number of sampling steps is $N$. Although many samplers are possible, in this paper we use the standard DDPM sampler [19].
Architecture
In early days, diffusion literature typically used UNets that consisted of ResNet blocks, with optional self-attention layers. More recent architecture either use full Transformers (DiT [35], StableDiffusion) or transformer backbones UViTs [21]. This transformer backbone makes it especially easy to use conditioning signals in these architecture via cross-attention layers. To be precise, one can say that the denoising neural network takes in ${\bm{z}}_{t}\in\mathbb{R}^{H\times W\times 3},t\in\mathbb{R},y\in\mathbb{R}^{T_{y}\times D_{y}}$. In principle it does not matter which diffusion model we use to generate images. In fact, every generative model would suffice. In practice we use the simple diffusion framework [21] because it can learn to generate high resolution images end-to-end without the need of a separate autoencoder.
4 Method
4.1 Episodic Dataset
The training dataset for our model is a subset of WebLI [9] that consists of roughly one billion images from the web. In particular, we use Episodic WebLI [8], where each episode contains randomly sampled loosely related images (i.e., they are clustered according to their URL). We randomly sample an image as the conditioning input and another image from the same episode as the ground-truth "target" image. We do not use any text inputs and only rely on image conditioning.
4.2 Image Encoder
Training an image-conditioned diffusion model requires two components: 1) An image encoder that extracts semantic information from a conditioning image. 2) A diffusion model that denoises a noisy target image conditioned on the image representations. We could train a separate Vision Transformer or ResNet model end-to-end as an image encoder with the diffusion model to learn useful conditioning representations.
Instead, we leverage pre-trained image encoders and condition our diffusion model on their “frozen" representations. This offers two advantages. First, we can precompute representations for all images in the dataset that eliminates expensive forward passes through the encoder during training. Second, by relying on pretrained encoders, we can just focus on scaling the diffusion model.
We investigate ViT image encoders trained with two pretraining strategies: training on web-text image pairs [37] (we use the SigLIP [55]), and self-supervision with teacher pseudo-labels (DINO) [6, 34]. Furthermore, we explore two conditioning strategies to incorporate conditioning information:
  1. 1.
    Token level conditioning: The diffusion model cross-attends to a sequence of tokens from the final transformer block of the image encoder.
  2. 2.
    Global feature conditioning: The encoder provides a single global feature representation from the encoder, being the CLS token for DINO and the output of the MAP block for SigLIP. We normalize and embed this with a dense projection and add it to the noise embedding. This subsequently conditions the diffusion model via FiLM layers [36] as done in standard label-conditioned diffusion.
4.3 Data Filtering
Each episode consists of images that are loosely related, whereas our model assumes conditioning and target images share semantic information. This mismatch may lead the model to waste capacity on modeling irrelevant noisy conditioning-target pairs. To address this, we filter out pairs with low semantic similarity. The pretrained encoder (DINO/SigLIP) computes the global CLS representation from the conditioning and target image. We compute the cosine similarity between the global conditioning/target representations and filter out pairs below a lower threshold. Additionally, to ensure generation of interesting images, we also filter conditioning-target pairs with similarity above a high threshold. This may ensure that generated images retain the semantics of the conditioning image, while being sufficiently different.
5 Experiments
5.1 Evaluation Datasets
We use $10\%$ of the Episodic WebLI dataset for the architecture design ablations and we use the entire dataset for the final model. We evaluate the model by conditioning it on ImageNet validation set images and then computing the FID between the generated samples and the ImageNet validation set. We investigate the transfer of Semantica on four datasets: ImageNet [39], LSUN-Outdoor Church, LSUN-Bedroom [54] and SUN-397 [50]. We report the FID for each dataset using 50K conditioning images randomly sampled from within the dataset. We use the validation dataset for ImageNet while the train set for all other datasets. To assess diversity, we compute the mean LPIPS [57] score between 50K generated samples and their corresponding conditioning images.
5.2 Image Encoder
Here, we evaluate the different encoder and conditioning design choices (Sec. 4.2). We train baseline diffusion models with two frozen image encoders (DINO-v2 B/14 and SigLIP B/16) and two conditioning strategies (cross-attention and FILM).
Refer to captionRefer to captionRefer to caption
Figure 2: We plot ImageNet FID as a function of number of training steps on Episodic WebLI. Left: FID of two image encoders DINO-v2 B/14 and SigLIP B/16 using two different conditioning methods: 1) All encoder output tokens via cross-attention (CA) 2) Global feature representation via FILM. Center: FID of DINO-v2 B/14 + Cross Attention with and without data filtering. Right: FID of DINO-v2 B/14 + Cross Attention with 1x1 patching and higher learning rate.
Fig. 2 left shows ImageNet FID as a function of number of training steps on Episodic WebLI. Token-level conditioning significantly outperforms global feature conditioning for both DINO-v2 and SigLIP, highlighting the importance of low-level features for this particular task. While SigLIP FiLM performs better than DINO-v2 FiLM, this trend reverses with cross-attention. This suggests a greater compatibility of DINO-v2 low-level features over CLIP low-level features for semantics-based image generation.
For this reason, we will now use DINO-v2 frozen representations with a cross-attention conditioning mechanism as our main architecture.
5.3 Data Filtering
We investigate the impact of semantic data filtering, as described in Sec. 4.3. We heuristically set the lower and higher thresholds to 0.3 and 0.9 respectively. Fig. 2 middle shows the FID of the DINO-v2 B/14 + Cross Attention with and without data filtering. Similarity-based data filtering in DINO feature space positively impacts the generation quality and improves FID by greater than 10. In future work, we can explore tuning these thresholds for a desired quality-diversity tradeoff or even directly conditioning the diffusion model on the desired similarity with the conditioning image.
5.4 Hyperparameters
For our baseline diffusion model, we inherit all hyper-parameters from the ImageNet label conditioned diffusion model. The denoising model follows a U-ViT architecture that operates on $256\times 256$ images. The architecture consists of a initial $2\times 2$ patchification layer [11, 45] that downsamples the noised image by a factor of 2. The model has four downsampling stages, where each stage downsamples the feature maps by a factor of 2 at its output. The resolution of the lowest feature map is $16\times 16$. Transformer blocks operate at the stages with the two lowest resolutions $16\times 16$ and $32\times 32$ and convolutional blocks operate in the remaining stages. The four downsampling stages have 128, 256, 512 and 1024 channels. The first three stages have three blocks each and the last stage has sixteen blocks. The optimizer is Adam [24] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1e^{-12}$, batch size of 2048 and a learning rate of $1e^{-4}$. We also use Polyak averaging with a decay factor of $0.9999$. The diffusion loss parameters include v-prediction with loss in epsilon phase and a cosine adjusted schedule with a noise resolution of 32. We use the DDPM sampler with an interpolation of 0.2 (standard deviation is $\sigma_{t\to s}^{0.2}\sigma_{st}^{0.8}$) and no guidance for our ablations. Each ablation run utilizes 256 TPUv3 [15] chips around 2 days. However, the consistent ranking of different ablations throughout training can allow for a much shorter training schedule to identify the best model.
We finally perform a lightweight sweep of the diffusion model hyper-parameters. Notably, two minor modifications improve generation quality: 1) Replacing the 2x2 patching with no patching. The model consists of a $1\times 1$ convolution at the input and output. The downsampling proceeds with an additional stage consisting of two convolutional blocks and 128 channels each. 2) Increasing the learning rate from $41e-$ 4 to $42e-$ 4. Fig. 2 right shows the improvement in FID of the DINO B/14 encoder with these modifications. Other changes such as tuning noise schedules, dropout rates or Polyak decay factors cause negligible changes or decrease sample quality.
5.5 Final Training
In our final training phase, we make the following four adjustments:
  1. 1.
    Replace DINO-v2 ViT-B/14 conditioning features with the larger DINO ViT-L/14 conditioning features.
  2. 2.
    Scale the model by doubling the number of channels across all UViT stages. The only exception is the compute-heavy first downsampling stage which we keep at 128.
  3. 3.
    Expand training data to include the entirety of episodic WebLI which was previously $10\%$ in the ablations.
  4. 4.
    Extend the training duration from 300K steps to two million steps
We train our model on 256 Cloud TPUv5 chips [14] for approximately 12 days with a constant learning rate of $2\text{e}{-4}$.
5.6 Label Grouped Baseline
Here we compare Semantica to a baseline that has direct label supervision (for example, ImageNet). Recall that the conditioning and target image belong to the same webpage. However, in the presence of label supervision (as in ImageNet), the target and conditioning image can just belong to the same class label. So as a supervised baseline, we group images on ImageNet as per their label and train Semantica on this dataset. Table 1 compares the FID of the Label Grouped baseline (LG) to Semantica. Since LG is trained on ImageNet, it significantly outperforms Semantica (FID 4.8 vs FID 18.4). However, this trend reverses on all other datasets, where Semantica outperforms LG. Both the supervised baseline and Semantica rely on the DINOv2 encoder which was trained on a wide variety of data sources. Therefore the encoder itself may provide useful representations on a number of datasets. But training LG just on ImageNet, might limit the diffusion model’s exposure from non ImageNet images, potentially explaining its significant performance drop on all other datasets.
5.7 One Shot Results
We evaluate Semantica’s one-shot image generation capabilities. Each row in Fig. 1 showcases a conditioning image to the left and five generated samples to the right. We observe that Semantica is capable of generating high quality and diverse images.
Table 1 reports quantitative results. Semantica receives 1000 and 397 conditioning images for ImageNet and SUN397 respectively, one per class. The model generates 50 and 120 samples per class for ImageNet and SUN397 totalling 50000 and 47840 generated samples. We then report the FID with 50000 ground truth samples. Semantica achieves a FID of 19.1 and 9.5 on ImageNet and SUN397 respectively. On increasing the guidance to 0.5, the FID improves to 9.3 and 7.1 respectively. Simply replicating the conditioning image Copy 50 and 120 times, achieves a much worse FID of 38.4 and 70.6.
5.8 Effect of Guidance
We investigate the quality/diversity tradeoff offered by Semantica’s guidance parameter. Fig. 3 Left illustrates the FID and LPIPS on ImageNet for different guidance factors. Note that we filter out near duplicates with a semantic similarity greater than 0.9 from the training set. This makes it unlikely that the model copies the conditioning image exactly to achieve a perfect FID. With no guidance, Semantica achieves a FID of 18.4 on ImageNet. The high LPIPS score 0.72 further reflects the model’s ability to generate diverse samples. On increasing the guidance upto a certain extent, we tradeoff quality with diversity. Notably, at guidance factor 0.5, which is commonly used in label conditioned generation, Semantica achieves a FID of 6.2 with a LPIPS score of 0.68. Fig. 4, displays five conditioning images from ImageNet and the generated samples at different guidance factors. At guidance factor 0.0, the samples reflect a broad semantic category from the conditioning image. Increasing the guidance factor leads to samples that incorporate more specific details from the conditioning image. For example, with the conditioning image of the dog and the kid, Semantica stars with a sample of a dog. The specific breed of the dog and the child in the image appear as we amplify the guidance.
Refer to captionRefer to caption
Figure 3: LPIPS (Higher is more diverse) vs FID (Lower is better) for five different datasets (Left: ImageNet, Right: LSUN Bedroom, LSUN Churches, SUN397). Each point represents a different guidance factor. As expected, we tradeoff quality with diversity on increasing guidance, but even with lower diversity the model does not collapse completely to its conditioning image.
ImageNet Bedroom Church SUN397
Label grouped 4.8 46.2 27.1 29.7
Semantica 18.4 6.2 17.3 6.7
Guidance @@@@ 0.5
Label grouped 5.1 34.2 20.4 22.4
Semantica 6.2 2.4 4.0 2.5
ImageNet SUN397
Semantica 19.1 9.5
+ Guidance @@@@ 0.5 9.3 7.1
Copy 38.4 70.6
Table 1: Left: Comparison between Semantica and a Label Grouped baseline (LG) trained on ImageNet. The conditioning and target pairs have the same label for LG. LG outperforms Semantica in-distribution and performs worse on out-of-distribution datasets. Right (One Shot Generation Results): Evaluates Semantica one-shot image generation capabilities. Semantica receives 1000 and 397 conditioning images for ImageNet and SUN397 respectively, one-per-class. The model generates 50 and 120 samples per-class for ImageNet and SUN397 totalling 50000 and 47840 generated samples. We then report the FID with 50000 ground truth samples. Copy is a baseline where the same per-class conditioning image is just replicated 50 and 120 times respectively.
Figure 4: Left: Conditioning Image from ImageNet. Right: Generated samples with guidance factors 0.0, 0.1, 0.2, 0.5 and 1.0. At guidance factor 0.0, the samples reflect a broad semantic category from the conditioning image. Increasing the guidance factor leads to samples that incorporate more specific details from the conditioning image.
Fig. 3 Right demonstrates quantitative results on three smaller datasets LSUN Bedroom, LSUN Church and SUN-397. We observe a similar tradeoff between quality and diversity as we vary the guidance factor. With zero guidance, Semantica achieves a FID of 6.2, 17.3 and 6.7 on LSUN Bedroom, LSUN Church and SUN397 respectively. On increasing the guidance to 0.5, the FID improves to 2.4, 4.0 and 2.5 respectively.
Fig. 5 showcases samples for each small dataset across various guidance factors. In row four, the bed (main object) persists across all guidance levels, while the chair and the fence appear at high guidance levels. Row five exhibits a similar effect: the number of minarets in the generated church increases from one to two and the shape of the main dome begins to resemble the conditioning image. In row one, the sample resembles toys with zero guidance, the sample resembles toys but transforms into a crowded convention as guidance increases.
SUN397
Refer to caption Refer to caption
Refer to caption Refer to caption
LSUN Bedrooms
Refer to caption Refer to caption
Refer to caption Refer to caption
LSUN Churches
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Left: Conditioning Images from SUN397 (Top two rows), LSUN Bedrooms (Middle two rows) and LSUN churches (Last two rows). Right: Generated samples with guidance factors 0.0, 0.1, 0.2, 0.5, 1.0 and 1.5.
6 Conclusion
We present Semantica, an image-conditioned diffusion model capable of generating high-quality and diverse images based on the semantics of a conditioning image. Our pretraining strategy is simple, the diffusion model models a target image with the conditioning image coming from the same webpage. Our experiments also further emphasize the expressivity of pretrained image encoders and necessity of semantic-based data filtering.
Limitations
We identify the following limitations of Semantica. 1) While adapting Semantica to downstream datasets, requires no finetuneing, a limitation is that training Semantica requires high computational resources as is the case with most generative models. However, any improvements that make diffusion models more efficient is likely to transfer to Semantica. 2) Our model depends on a frozen encoder and we do not modify the weights. With more resources, we can also study the tradeoffs with this design decision. 3) With a very high guidance factor (>1.0) , Semantica might converge to an oversaturated conditioning image. An area of improvement would be to find ways to steer it to another image with the same semantic information. 4) Our model as is does not incorporate additional conditioning signals. Generalizing the model to utilize additional conditioning information can be an exciting area of future research. 5) While our model generally produces diverse and high-quality images, sometimes artifacts can appear especially at guidance 0.0 and in areas like lower-level structure and human faces. Scaling our model further may remove these artifacts.
Appendix A Broader Impact Statement
This paper introduces a model capable of generating images that reflect the semantic information of a given input image. Potential positive applications include being a creative aid, augmentation of an image retrieval system or enhancing the size of an image dataset. While the model lacks direct controllability thus mitigating some negative applications like creation of personalized deepfakes, our model can still generate deepfakes. Therefore, we advocate for responsible usage for our model.
Appendix B Standard Errors
In general, the variance in FID is small. We compute the variance across different random seeds and different randomly sampled one shot datasets on ImageNet. With a fixed one-shot dataset and five different random seeds, we report a FID of $9.4\pm 0.02$. We also sample five different one-shot datasets and report a FID of $9.5\pm 0.06$.
Appendix C ImageNet Samples
Figure 6: Set 1: Conditioning Images from ImageNet
Figure 7: Set 1: Random Samples 1
Figure 8: Set 1: Random Samples 2
Figure 9: Set 1: Random Samples 3
Figure 10: Set 2: Conditioning Images from ImageNet
Figure 11: Set 2: Random Samples 1
Figure 12: Set 2: Random Samples 2
Figure 13: Set 2: Random Samples 3
Figure 14: Set 3: Conditioning Images from ImageNet
Figure 15: Set 3: Random Samples 1
Figure 16: Set 3: Random Samples 2
Figure 17: Set 3: Random Samples 3
Appendix D LSUN Bedroom Samples
Figure 18: Conditioning Images from LSUN Bedroom
Figure 19: Random Samples 1
Figure 20: Random Samples 2
Figure 21: Random Samples 3
Appendix E SUN 397 Samples
Figure 22: Conditioning Images from SUN397
Figure 23: Random Samples 1
Figure 24: Random Samples 2
Figure 25: Random Samples 3
Appendix F LSUN Church Samples
Figure 26: Set 1: Conditioning Images from LSUN Churches
Figure 27: Random Samples 1
Figure 28: Random Samples 2
Figure 29: Random Samples 3